At one instant an electron(charge = -1.6 x 10-19 C)is moving in the xy plane , the components of its velocity beingvx = 3.0 x 105 m/s and vy=4.0x105m/s. A magnetic field of 0.8T is in the positivez direction. At that instant the magnitude of the magnetic force onthe electron is ....
(a) 0
(b) 2.9 x 10-14 N
(c) 3.8 x 10-14 N
(d) 4.8 x 10-14 N
(e) 6.4x10-14 N
Answer #1
v =vxi+vyj =3.0 x 105i +4.0x105j m/s

B = 0.8 k

q =
-1.6 x 10-19 C

F = q(v X B)=
-1.6 x10-19(3.0 x105i + 4.0x105j) X 0.8 k
=
-1.6 x 10-19 *3.0 x 105 *0.8(-j) + -1.6 x10-19 *4.0x105*0.8(-i)
= 3.84 x 10-14j + 5.12 x10-14i [ Sincei X k = -j andj X k = -i]

∴ |F| = √[( 3.84 x 10-14)2 +(5.12 x10-14)2] = 6.4 x10-14 N